40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 19 09:36:32 2013, Koji, Update, Green Locking, Xend green aligned ALS_OUTOFLOOP_130816.pdf
    Reply  Mon Aug 19 14:22:36 2013, rana, Update, Green Locking, Xend green aligned 
       Reply  Mon Aug 19 17:40:32 2013, Steve, Update, Green Locking, Xend green layout corrections beforeC.jpgnowC.jpgstillMultiple.jpg
          Reply  Mon Aug 19 19:08:35 2013, Koji, Update, Green Locking, Xend green layout corrections 
          Reply  Tue Aug 20 10:59:15 2013, Steve, Update, Green Locking, Xend green layout corrections faraway.jpgcloser.jpg
             Reply  Wed Aug 28 11:15:02 2013, Steve, Update, Green Locking, Xend green layout corrections 
    Reply  Mon Aug 19 16:18:56 2013, manasa, Update, Green Locking, Xend green aligned 
Message ID: 9035     Entry time: Mon Aug 19 19:08:35 2013     In reply to: 9034
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: Xend green layout corrections 

- An Aluminum mirror instead of 2" unknown mirror for the pick-off for the rejected beam from the green faraday isolator (Steve)
=> Replaced. To be reviewed

- Faraday mount replacement. Check what we have for the replacement. (Steve)

- The green REFL PD should be closer to the pick-off mirror. (Steve)
=> Moved. To be reviewed

- A beam dump should be placed for the green REFL PD

- Move the green shutter to the place where the spot is small (Steve)
=> Moved. To be reviewed.

- The pole of the PZT mounting should be replaced with a reasonable one. (Steve with Manasa's supervision)

- Tidying up doubling oven cable. Make a hole on the wall. (Steve)
=> Done. To be reviewed.

- Tidying up the PZT cabling (Steve)

- The optics are dirty. To be drag wiped. (Manasa, Masayuki)

ELOG V3.1.3-